OpenStudy (anonymous):

if f(x)=tanx-x and g(x)=x^3, evaluate the limit of f(x) over g(x) as x approaches 0. -3rd question.

OpenStudy (anonymous):

@zarkon, et al, apparently she needs a non-l'hopital method, and i'll be darned if i can think of one

OpenStudy (zarkon):

how about a power series solution :)

OpenStudy (anonymous):

yes, do not pass go, do not collect $200

OpenStudy (zarkon):

lol

OpenStudy (zarkon):

go directly to jail

OpenStudy (anonymous):

in fact i think you need l'hopital twice at least

OpenStudy (zarkon):

power series gives the quickest solution

OpenStudy (anonymous):

\[\frac{\sec^2(x)-1}{3x^2}\] \[\frac{\tan^2(x)}{3x^2}\] etc

OpenStudy (anonymous):

yes but if this is pre-derivative, then what can you do?

OpenStudy (nenadmatematika):

kaezalorene you can always say to your teacher that you've discovered some rule when you were home studying.....the rule that you can use the derivatives and solve....he will be suprised :D

OpenStudy (nenadmatematika):

kidding :D

OpenStudy (ash2326):

\[f(x)=\frac{sin x}{cosx}-x\] \[g(x)=x^3\] we have \[f(x)/ g(x)=\frac{\frac{sin x}{cosx}-x}{x^3}\] dividing numerator and denominator by x now \[f(x)/g(x)= \frac{\frac{sin x}{x}-cos x}{x^2}\] \[limit x-->0 \frac{sinx}{x}=1\] so we have \[f(x)/g(x)= \frac{1-cos x}{x^2}\] 1-cos x = 2 (sin x/2)^2 so we get \[f(x)/g(x)= \frac{2 sin ^2 x/2}{x^2}\] dividing the numerator and denominator by 4 we get \[f(x)/g(x)=\frac {\frac{2 sin ^2 x/2}{4}}{\frac{x^2}{4}}\] \[(sin x / 2)/ (x/2-)-->1 as x---->0\] so we get \[f(x)/g(x)= \frac {2}{4}\] = 1/2

OpenStudy (ash2326):

I missed a cos x in the denominator but it won't make any difference as x-->0 cos x ---> 1

OpenStudy (ash2326):

Zarkon what do you think , is my method correct?

OpenStudy (zarkon):

the answer is 1/3

OpenStudy (zarkon):

unfortunatly you can't do it like that (the sin(x)/x is an integral part of the final solution...you can't takes its limit while holding everything else constant)

OpenStudy (ash2326):

Thanks I got your point